# Path-dependent Hamilton–Jacobi equations with $u$-dependence and time-measurable Hamiltonians

### v1

@@ Abstract.
We establish existence and uniqueness of minimax solutions for a fairly general class of path-dependent Hamilton–Jacobi equations. In particular, the relevant Hamiltonians can contain the solution and they only need to be measurable with respect to time. We apply our results to optimal control problems of (delay) functional differential equations with cost functionals that have discount factors and with time-measurable data. Our main results are also crucial for our companion paper Bandini and Keller (2024), where non-local path-dependent Hamilton–Jacobi–Bellman equations associated to the stochastic optimal control of non-Markovian piecewise deterministic processes are studied.
@@
The research of the first named author was partially supported by the 2018 GNAMPA-INdAM project Controllo ottimo stocastico con osservazione parziale: metodo di randomizzazione ed equazioni di Hamilton-Jacobi-Bellman sullo spazio di Wasserstein and the 2024 GNAMPA-INdAM Problemi di controllo ottimo in dimensione infinita. The research of the second named author was supported in part by NSF-grant DMS-2106077.
@@ 1. Introduction
with time-measurable Hamiltonian, i.e., the mappings $t\mapsto H(t,x,y,z)$ only need to be Borel-measurable. The path space $D([0,T],{\mathbb{R}}^{d})$ in (1.1) is the set of all right-continuous functions from $[0,T]$ to ${\mathbb{R}}^{d}$ that have left limits. Important special cases of (1.1) are Isaacs equations associated to differential games with history-dependent data and Hamilton–Jacobi–Bellman (HJB) equations associated to optimal controls problems involving (delay) functional differential equations. There exists already a large body of literature for non-smooth solutions of those equations with continuous Hamiltonians (see [GLP21AMO] and the references therein). Typically, $C([0,T],{\mathbb{R}}^{d})$ is used as path space. But this difference is not material.
@@
The two most common notions of non-smooth solutions for path-dependent Hamilton–Jacobi equations are viscosity solutions and minimax solutions. In fact, for a large class of path-dependent equations but with continuous (!) Hamiltonians, these two notions turn out to be equivalent (see [GP23JFA]).111 Note that we do not pursue here a generalization of those equivalence results in our setting.
@@
Here, we work with minimax solutions. This helps a lot in our setting, where $H$ in (1.1) is time-measurable, as no change of notion is needed compared to the case of $H$ being continuous. Furthermore, we also extend the theory of minimax solutions for path-dependent equations in [Lukoyanov03] to the case of $u$-dependent Hamiltonians and thereby provide a counterpart to the theory of minimax solutions for (non-path-dependent) Hamilton–Jacobi equations with $u$-dependent Hamiltonians in [Subbotin] (see also the references therein for earlier works but with more restrictive assumptions).
@@
Viscosity solutions still play a useful role. They are utilized for the proof of our comparison principle.222 We use the same methodology for establishing existence and uniqueness for minimax solutions with the help of viscosity solution techniques as in [BK18JFA]: The “five-step-scheme” described on p. 2103 therein to be more precise.
@@ 1.1. Further related literature
Results for viscosity solutions for (non-path-dependent) equations with only time-measurable Hamiltonians were first published in [Ishii85] (this work had also some influence for the proof of our comparison principle in section 5). Further related early works are [LionsPerthame87] and [BarronJensen87].
Minimax solutions for (non-path-dependent) equations with time-measurable Hamiltonians were studied in [SonEtAl95, SonEtAl97, minimaxBookTimeMeas, VanEtAl95]. Note that these works unlike ours require positive homogeneity of the Hamiltonian with respect to the gradient.
@@
1.2. A difficulty and its resolution
@@ 1.3. Organization of the rest of the paper
Section 2 introduces the setting (path space with topology) and some notation. In section 3, we give meaning to the derivatives $\partial_{t}u$ and $\partial_{x}u$ in (1.1). Section 4 contains the definition of minimax solution to terminal-value problems involving (1.1) as well as standing assumptions for our data such as the Hamiltonian $H$. In section 5, a comparison principle for (1.1) is established. In section 6, we establish a general existence result via Perron’s method. Finally, in section 7, we consider an optimal control problem for (delay) functional differential equations with time-measurable data and we show that the value function is the unique minimax solution to the associated HJB equation.
@@ 2. Setting and notation
$\mathbf{d}_{\infty}((t,x),(s,\tilde{x})):=\left|{t-s}\right|+\sup_{0\leq r\leq T}\left|{x(r\wedge t)-\tilde{x}(r\wedge s)}\right|.$
@@
$\mathcal{X}^{L}(s_{0},x_{0})$ :=$:=\{x\in\Omega:\,x=x_{0}\text{ on $[0,s_{0}]$, $x|_{[s_{0},T]}$ is absolutely continuous with}$
$\qquad\qquad\left|{x^{\prime}(t)}\right|\leq L(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ a.e.~{}on $(t_{0},T)$}\}.$
Those sets of “Lipschitz-like” paths are very important. In particular, thanks to their useful properties listed in the following remarks, they are helpful insofar they circumvent difficulties coming from the lack of local compactness of $\Omega$.
@@ 3. Path derivatives
We write $\varphi\in C^{1,1,1}([0,T]\times\Omega\times\Omega)$ if $\varphi\in C([0,T]\times\Omega\times\Omega)$ and there are functions $\partial_{t}\varphi\in C([0,T]\times\Omega\times\Omega)$ and $\partial_{x}\varphi$, $\partial_{\tilde{x}}\varphi\in C([0,T]\times\Omega\times\Omega,{\mathbb{R}}^{d})$, called path derivatives of $\varphi$, such that, for every $(t_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, every pair $(x,\tilde{x})\in C([0,T]\times\Omega)$ with $(x,\tilde{x})|_{[0,t_{0}]}=(x_{0},\tilde{x}_{0})|_{[0,t_{0}]}$ and $(x,\tilde{x})|_{[t_{0},T]}$ being Lipschitz continuous, and every $t\in(t_{0},T]$, we have
$\varphi(t,x,\tilde{x})-\varphi(t_{0},x_{0},\tilde{x}_{0})$
$\qquad=\int_{t_{0}}^{t}\left[\partial_{t}\varphi(s,x,\tilde{x})+x^{\prime}(s)\cdot\partial_{x}\varphi(s,x,\tilde{x})+\tilde{x}^{\prime}(s)\cdot\partial_{\tilde{x}}\varphi(s,x,\tilde{x})\right]\,{\mathrm{d}}s.$
@@ 4. Minimax solutions
Next, we introduce sets of paths needed in our definition of minimax solution.555 Note that without the $u$-dependence of $H$, the situation would be much easier. Only the sets $\mathcal{X}^{L}(s_{0},x_{0})$ (and not $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$) would then be needed (see, e.g., [BK18JFA, section 1.3]).
@@
$\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ :=$:=\{(x,y)\in\mathcal{X}^{L}(s_{0},x_{0})\times C([t_{0},T]):$
$\qquad\qquad y(t)=y_{0}+\int_{s_{0}}^{t}\left[(x^{\prime}(s),z)-H(s,x,y(s),z)\right]\,{\mathrm{d}}s\text{ on $[t_{0},T]$}\}.$
@@
Thanks to Assumption 4.2 (i) and (vi), the sets $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ are non-empty (cf. Proposition 2.13 (i) in [BK18JFA]) and compact in $(\Omega\times C([t_{0},T]),\lVert{\cdot}\rVert_{\infty})$ (cf.  Remark 2.1). Moreover, using additionally Assumption 4.2 (iii), one can show that also the intersections $\mathcal{Y}^{L_{H}}(s_{0},x_{0},y_{0},z)\cap\mathcal{Y}^{L_{H}}(s_{0},x_{0},y_{0},\tilde{z})$, $z\neq\tilde{z}$, are non-empty (see, e.g., [Subbotin, pages 73-74] or [BK18JFA, pages 2124-2125]).
@@
Notice that classical solutions are minimax solutions. We sketch some of the arguments (adapted from [Subbotin, section 2.4 ]). Assume that $H$ is continuous and that $u$ is a classical solution of (4.1), i.e., $u\in C^{1,1}([0,T]\times\Omega)$ (this space is an obvious modification666 Just assume that $(t,x,\tilde{x})\mapsto u(t,x)\in C^{1,1,1}([0,T]\times\Omega)$. of Definition 3.1) and, for every $(t,x)\in[0,T)\times\Omega$,
@@
$x^{\prime}(t)=\begin{cases}\frac{H(t,x,u(t,x),\partial_{x}u(t,x))-H(t,x,u(t,x),z)}{\left|{\partial_{x}u(t,x)-z}\right|^{2}}\cdot(\partial_{x}u(t,x)-z)&\text{if $u(t,x)\neq z$,}\\ 0&\text{if $u(t,x)=z$}\end{cases}.$
@@
$y^{\prime}(t)=(x^{\prime}(t),z)-H(t,x,u(t,x),z)\quad\text{a.e.~{}on $(s_{0},T)$}$
@@
$\frac{d}{dt}u(t,x)$ $=\frac{d}{dt}\int_{s_{0}}^{t}\partial_{t}u(s,x)+(x^{\prime}(s),\partial_{x}u(s,x))\,ds$
$=-H(t,x,u(t,x),\partial_{x}u(t,x))+(x^{\prime}(t),\partial_{x}u(t,x)-z)+(x^{\prime}(t),z)$
$=-H(t,x,u(t,x),z)+(x^{\prime}(t),z)=y^{\prime}(t),$
@@
By compactness777 Note that Assumption 4.2 (vi) prevents a possible blow up of $y^{m}$ as $m\to\infty$. of $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ (Remark 4.3), we can, without loss of generality, assume that $(x^{m},y^{m})_{m}$ converges uniformly to a pair $(x^{0},y^{0})\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$. Let $t\in(s_{0},T]$. Then there is a sequence $(s^{m})_{m}$ in $(s_{0},T]$ with $s^{m}\in\pi^{m}$, $m\in{\mathbb{N}}$, that converges to $t$. Thus $y^{0}(t)=\lim_{m}y^{m}(s^{m})\geq\liminf_{m}u(s^{m},x^{m})\geq u(t,x^{0})$ thanks to the lower semi-continuity of $u$. This concludes the proof. ∎

### v2

@@ Abstract.
We establish existence and uniqueness of minimax solutions for a fairly general class of path-dependent Hamilton–Jacobi equations. In particular, the relevant Hamiltonians can contain the solution and they only need to be measurable with respect to time. We apply our results to optimal control problems of (delay) functional differential equations with cost functionals that have discount factors and with time-measurable data. Our main results are also crucial for our companion paper Bandini and Keller [arXiv preprint arXiv:2408.02147 (2024)], where non-local path-dependent Hamilton–Jacobi–Bellman equations associated to the stochastic optimal control of non-Markovian piecewise deterministic processes are studied.
The authors would like to thank the anonymous referees for many valuable comments.
@@ 1. Introduction
with time-measurable Hamiltonian, i.e., the mappings $t\mapsto H(t,x,y,z)$ only need to be Borel-measurable. The path space $D([0,T],{\mathbb{R}}^{d})$ in (1.1) is the set of all right-continuous functions from $[0,T]$ to ${\mathbb{R}}^{d}$ that have left limits. Important special cases of (1.1) are Isaacs equations associated to differential games with history-dependent data and Hamilton–Jacobi–Bellman (HJB) equations associated to optimal controls problems involving (delay) functional differential equations. Consider, e.g., a value function $v$ of the form
$v(t_{0},x_{0})=\inf_{a(\cdot)\in\mathcal{A}}h(\{x^{t_{0},x_{0},a(\cdot)}(t)\}_{0\leq t\leq T}),\quad(t_{0},x_{0})\in[0,T]\times D([0,T],{\mathbb{R}}^{d}),$
where $x=x^{t_{0},x_{0},a(\cdot)}:[0,T]\to{\mathbb{R}}^{d}$ solves an equation of the form
$x^{\prime}(t)$ $=f(t,x(t-1),a(t))\text{ on $(t_{0},T)$, where $x(s):=x(0)$ in case $s<0$,}$
$x|_{[0,t_{0}]}$ $=x_{0}|_{[0,t_{0}]}.$
Then $v$ should formally solve the HJB equation
$\partial_{t}v(t,x)+\inf_{a\in A}\left[f(t,x({\max\{t-1,0\}}),a)\cdot\partial_{x}v(t,x)\right]$ $=0,\,\,(t,x)\in[0,T)\times D([0,T],{\mathbb{R}}^{d}),$
$v(T,x)$ $=h(x),\quad x\in D([0,T],{\mathbb{R}}^{d}).$
[…]
(1.2) $x|_{[0,t]}=\tilde{x}|_{[0,t]}\Longrightarrow u(t,x)=u(t,\tilde{x})\text{ and }H(t,x,y,z)=H(t,\tilde{x},y,z),$
in other words $u(t,x)$ and $H(t,x,y,z)$ depend only on the history of the path $x$ until time $t$, i.e., on $\{x(s)\}_{0\leq s\leq t}$. This requirement makes equations of the form (1.1) as well as their 2nd order counterparts suitable for deterministic and stochastic optimal control problems with history dependent data (the controller knows the history and the current state but not the future). Second, the so-called path derivatives $\partial_{t}$ and $\partial_{x}$ in (1.1) (originated in [Kim85]) are not understood in the usual (Fréchet) sense. Instead they are defined in a way such that a relatively simple chain rule holds, e.g., $\frac{d}{dt}u(t,x)=\partial_{t}u(t,x)+\partial_{x}u(t,x)\cdot x^{\prime}(t)$ for a non-anticipating “smooth” function $u$ on $[0,T]\times D([0,T],{\mathbb{R}}^{d})$ and a differentiable path $x(\cdot)$ on $[0,T]$. Nevertheless, these path derivatives are compatible with the usual derivatives in the non-path-dependent case (see below in section 3 for details concerning path derivatives).
@@
Just as for “standard” Hamilton–Jacobi equations on finite-dimensional spaces, we do not expect to have smoothness, so we need non-smooth solutions. The two most common notions of non-smooth solutions for path-dependent Hamilton–Jacobi equations are viscosity solutions and minimax solutions.222 Minimax solutions originated in the theory of differential games (see, e.g., [KrasovksiiSubbotin]), where minmax operations are common, which motivates the name minimax solutions. See also Remark 4.6 below for more motivation. In fact, for a large class of path-dependent partial differential equations but with continuous (!) Hamiltonians, these two notions turn out to be equivalent (see [GP23JFA]).333 Note that we do not pursue here a generalization of those equivalence results in our setting.
Note that a large body of literature for non-smooth solutions of path-dependent partial differential equations with continuous Hamiltonians has been created (see [GLP21AMO, GL24survey] and the references therein). Among the early works, we want to mention [Lukoyanov2000, Lukoyanov03, Lukoyanov07, AubinHaddad02PPDE] in the 1st order case and [EKTZ11, ETZ_I, ETZ_II] in the 2nd order case. Also note that typically, $C([0,T],{\mathbb{R}}^{d})$ is used as path space. But this difference is not material.
@@
Here, we work with minimax solutions. This helps a lot in our setting, where $H$ in (1.1) is time-measurable, as no change of definition is needed compared to the case of $H$ being continuous. Furthermore, we also extend the theory of minimax solutions for path-dependent equations in [Lukoyanov03] to the case of $u$-dependent Hamiltonians and thereby provide a counterpart to the theory of minimax solutions for non-path-dependent Hamilton–Jacobi equations with $u$-dependent Hamiltonians in [Subbotin] (see also the references therein for earlier works but with more restrictive assumptions).
@@
Viscosity solutions still play a useful role. They are utilized for the proof of our comparison principle.444 We use the same methodology for establishing existence and uniqueness for minimax solutions with the help of viscosity solution techniques as in [BK18JFA]: The “five-step-scheme” described on p. 2103 therein to be more precise.
@@ 1.1. Further related literature
Results for viscosity solutions for non-path-dependent equations with only time-measurable Hamiltonians were first published in [Ishii85] (this work had also some influence for the proof of our comparison principle in section 5). Further related early works are [LionsPerthame87] and [BarronJensen87].
Minimax solutions for non-path-dependent equations with time-measurable Hamiltonians were studied in [SonEtAl95, SonEtAl97, minimaxBookTimeMeas, VanEtAl95]. Note that these works unlike ours require positive homogeneity of the Hamiltonian with respect to the gradient.
@@
Viscosity solutions for path-dependent Hamilton–Jacobi equations related to systems with “distributed time delays” only were investigated in [Lukoyanov07] and related to systems with “distributed time delays” as well as finitely many “discrete time delays” in [Lukoyanov10b]. A breakthrough in [zhou2020viscosity], where for the first time the smooth functional $\Psi$ (see (5.6) below) was studied, led to much more general results, in particular, existence and uniqueness for viscosity solutions of path-dependent HJB equations with Hamiltonians that only need to be Lipschitz continuous in the path variable $x=x(\cdot)$ with respect to the supremum norm $x\mapsto\sup_{t\in[0,T]}\left|{x(t)}\right|$. The mentioned functional $\Psi$ was crucial in [zhou2020viscosity] to prove the comparison principle for viscosity solutions and it is also crucial for our work (see Theorem 5.3 and its proof below).
1.2. Some difficulties and their resolutions
@@
We use Perron’s method to obtain our existence result. It should be noted that continuity of the data is usually used in applying Perron’s method for viscosity solutions (see [Ishii87Perron]) and we do not know how to overcome the lack of continuity of the Hamiltonian with respect to time, especially given that we work with partial differential equations on an infinite-dimensional space. Using minimax solutions circumvents this issue.
@@ 1.3. Organization of the rest of the paper
Section 2 introduces the setting (path space with topology) and some notation. In section 3, we give meaning to the derivatives $\partial_{t}u$ and $\partial_{x}u$ in (1.1). Section 4 contains the definition of minimax solution to terminal-value problems involving (1.1) as well as standing assumptions for our data such as the Hamiltonian $H$. In section 5, a comparison principle for (1.1) is established. In section 6, we establish a general existence result via Perron’s method. In section 7, we consider an optimal control problem for (delay) functional differential equations with time-measurable data and we show that the value function is the unique minimax solution to the associated path-dependent HJB equation. In section 8, we consider non-path-dependent counterparts of our previous results. In particular, we establish that the value function of an optimal control problem with time-measurable data is the unique minimax solution of a standard non-path-dependent HJB equation. Such a result is also new to the best of our knowledge. Finally, in the appendix, we establish regularity of the value function for our optimal control problem in section 7.
@@ 2. Setting and notation
$\mathbf{d}_{\infty}((t,x),(s,\tilde{x})):=\left|{t-s}\right|+\sup_{0\leq r\leq T}\left|{x(r\wedge t)-\tilde{x}(r\wedge s)}\right|,$
where $r\wedge t:=\min\{r,t\}$ for any $r$, $t\in{\mathbb{R}}$.
Remark 2.1.
Semicontinuous functions on $[0,T]\times\Omega$ equipped with $\mathbf{d}_{\infty}$ are automatically non-anticipating in the sense of (1.2).
@@
(2.1) $\begin{split}&\mathcal{X}^{L}(s_{0},x_{0}):=\Bigl{\{}x\in\Omega:\,\text{$x|_{[s_{0},T]}$ is absolutely continuous with}\\ &\qquad\qquad\left|{x^{\prime}(t)}\right|\leq L(1+\sup_{s\leq t}\left|{x(s)}\right|)\text{ a.e.~{}on $(s_{0},T)$ and $x=x_{0}$ on $[0,s_{0}]$}\Bigr{\}}.\end{split}$
Those sets of Lipschitz paths are very important. In particular, thanks to their useful properties listed in the following remarks, they are helpful insofar as they circumvent difficulties coming from the lack of local compactness of $\Omega$.
@@ 3. Path derivatives
We write $\varphi\in C^{1,1,1}([0,T]\times\Omega\times\Omega)$ if $\varphi\in C([0,T]\times\Omega\times\Omega)$ and there are functions $\partial_{t}\varphi\in C([0,T]\times\Omega\times\Omega)$ and $\partial_{x}\varphi$, $\partial_{\tilde{x}}\varphi\in C([0,T]\times\Omega\times\Omega,{\mathbb{R}}^{d})$, called path derivatives of $\varphi$, such that, for every $(t_{0},x_{0},\tilde{x}_{0})\in[0,T)\times\Omega\times\Omega$, every pair $(x,\tilde{x})\in{\color[rgb]{0,0,0}\Omega\times\Omega}$ with $(x,\tilde{x})|_{[0,t_{0}]}=(x_{0},\tilde{x}_{0})|_{[0,t_{0}]}$ and $(x,\tilde{x})|_{[t_{0},T]}$ being Lipschitz continuous, and every $t\in(t_{0},T]$, we have
(3.1) $\begin{split}&\varphi(t,x,\tilde{x})-\varphi(t_{0},x_{0},\tilde{x}_{0})\\ &\qquad=\int_{t_{0}}^{t}\left[\partial_{t}\varphi(s,x,\tilde{x})+x^{\prime}(s)\cdot\partial_{x}\varphi(s,x,\tilde{x})+\tilde{x}^{\prime}(s)\cdot\partial_{\tilde{x}}\varphi(s,x,\tilde{x})\right]\,{\mathrm{d}}s.\end{split}$
@@
Independently, B. Dupire introduced in [dupirefunctional] explicitly defined path derivatives on $D([0,T],{\mathbb{R}}^{d})$ and established a functional Itô calculus. [^7] I.e., a non-Markovian extension of the usual Itô calculus. Dupire’s derivatives $\partial^{\mathrm{Dupire}}_{t}$, $\partial^{\mathrm{Dupire}}_{x}=(\partial^{\mathrm{Dupire}}_{x^{1}},\ldots,\partial^{\mathrm{Dupire}}_{x^{d}})$, and $\partial^{\mathrm{Dupire}}_{\tilde{x}}=(\partial^{\mathrm{Dupire}}_{\tilde{x}^{1}},\ldots,\partial^{\mathrm{Dupire}}_{\tilde{x}^{d}})$ for a function $\varphi:[0,T]\times\Omega\times\Omega\to{\mathbb{R}}$ are defined as follows (provided the limits below exist and are finite):
(3.2) $\small\begin{split}\partial^{\mathrm{Dupire}}_{t}\varphi(t,x,\tilde{x})&:=\lim_{\delta\downarrow 0}\frac{\varphi(t+\delta,x(\cdot\wedge\delta),\tilde{x}(\cdot\wedge\delta))-\varphi(t,x,\tilde{x})}{\delta}\text{ (provided $t<T$),}\\ \partial^{\mathrm{Dupire}}_{t}\varphi(T,x,\tilde{x})&:=\lim_{t\uparrow T}\partial^{\mathrm{Dupire}}_{t}\varphi(t,x,\tilde{x}),\\ \partial^{\mathrm{Dupire}}_{x^{i}}\varphi(t,x,\tilde{x})&:=\lim_{r\to 0}\frac{\varphi(t,(x^{1},\ldots,x^{i-1},x^{i}+r{\mathbf{1}}_{[t,T]},x^{i+1},\ldots,x^{d}),\tilde{x})-\varphi(t,x,\tilde{x})}{\delta},\\ \partial^{\mathrm{Dupire}}_{\tilde{x}^{i}}\varphi(t,x,\tilde{x})&:=\lim_{r\to 0}\frac{\varphi(t,x,(\tilde{x}^{1},\ldots,\tilde{x}^{i-1},\tilde{x}^{i}+r{\mathbf{1}}_{[t,T]},\tilde{x}^{i+1},\ldots,\tilde{x}^{d}))-\varphi(t,x,\tilde{x})}{\delta}\end{split}$
for each $i\in\{1,\ldots,d\}$. These derivatives coincide with the path derivatives in Definition 3.1 in the following sense: Suppose that $\varphi\in C([0,T]\times\Omega\times\Omega)$ is continuously differentiable in the sense of Dupire, i.e, $(\partial^{\mathrm{Dupire}}_{t}\varphi,\partial^{\mathrm{Dupire}}_{x}\varphi,\partial^{\mathrm{Dupire}}_{\tilde{x}}\varphi)$ defined by (3.2) exists and belong to $C([0,T]\times\Omega\times\Omega,{\mathbb{R}}\times{\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$. Then [zhou2020viscosity, Theorem 2.4] yields the chain rule (3.1) with $(\partial_{t}\varphi,\partial_{x}\varphi,\partial_{\tilde{x}}\varphi)$ replaced by $(\partial^{\mathrm{Dupire}}_{t}\varphi,\partial^{\mathrm{Dupire}}_{x}\varphi,\partial^{\mathrm{Dupire}}_{\tilde{x}}\varphi)$. It follows immediately that $\varphi\in C^{1,1,1}([0,T]\times\Omega\times\Omega)$ and, together with Remark 3.2, we can deduce that
$\partial_{t}\varphi=\partial^{\mathrm{Dupire}}_{t}\varphi,\quad\partial_{x}\varphi=\partial^{\mathrm{Dupire}}_{x}\varphi,\quad\partial_{\tilde{x}}\varphi=\partial^{\mathrm{Dupire}}_{\tilde{x}}\varphi.$
@@ 4. Minimax solutions
Next, we introduce sets of paths needed in our definition of minimax solution.888 Note that without the $u$-dependence of $H$, the situation would be much easier. Only the sets $\mathcal{X}^{L}(s_{0},x_{0})$ (and not $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$) would then be needed (see, e.g., [BK18JFA, section 1.3]).
@@
$\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ :=$:=\Bigl{\{}(x,y)\in\mathcal{X}^{L}(s_{0},x_{0})\times C([s_{0},T]):$
$\qquad\qquad y(t)=y_{0}+\int_{s_{0}}^{t}\left[{\color[rgb]{0,0,0}x^{\prime}(s)\cdot z}-H(s,x,y(s),z)\right]\,{\mathrm{d}}s\text{ on $[s_{0},T]$}\Bigr{\}}.$
@@
Thanks to Assumption 4.2 (i) and (vi), the sets $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ are non-empty (cf. Proposition 2.13 (i) in [BK18JFA]) and compact in $(\Omega\times C([s_{0},T]),\lVert{\cdot}\rVert_{\infty})$ (cf.  Remark 2.2). Moreover, using additionally Assumption 4.2 (iii), one can show that also the intersections $\mathcal{Y}^{L_{H}}(s_{0},x_{0},y_{0},z)\cap\mathcal{Y}^{L_{H}}(s_{0},x_{0},y_{0},\tilde{z})$, $z\neq\tilde{z}$, are non-empty (see, e.g., [Subbotin, pages 73-74] or [BK18JFA, pages 2124-2125]).
In the next remark, we use $D([0,T])$, the set of all right-continuous functions from $[0,T]$ to ${\mathbb{R}}$ that have left limits.
@@
Notice that classical solutions are minimax solutions. We sketch some of the arguments (adapted from [Subbotin, section 2.4 ]). Assume that $H$ is continuous and that $u$ is a classical solution of (4.1), i.e., $u\in C^{1,1}([0,T]\times\Omega)$ (this space is an obvious modification999 Just assume that $(t,x,\tilde{x})\mapsto u(t,x)\in C^{1,1,1}([0,T]\times\Omega)$. of Definition 3.1) and, for every $(t,x)\in[0,T)\times\Omega$,
@@
$x^{\prime}(t)=\begin{cases}\frac{H(t,x,u(t,x),\partial_{x}u(t,x))-H(t,x,u(t,x),z)}{\left|{\partial_{x}u(t,x)-z}\right|^{2}}\cdot(\partial_{x}u(t,x)-z)&\text{if $\partial_{x}u(t,x)\neq z$,}\\ 0&\text{if $\partial_{x}u(t,x)=z$}\end{cases}$
@@
$y^{\prime}(t)={\color[rgb]{0,0,0}x^{\prime}(t)\cdot z}-H(t,x,u(t,x),z)\quad\text{a.e.~{}on $(s_{0},T)$}$
@@
$\frac{d}{dt}u(t,x)$ $=\frac{d}{dt}\int_{s_{0}}^{t}\partial_{t}u(s,x)+{\color[rgb]{0,0,0}x^{\prime}(s)\cdot\partial_{x}u(s,x)}\,ds$
$=-H(t,x,u(t,x),\partial_{x}u(t,x))+{\color[rgb]{0,0,0}x^{\prime}(t)\cdot[\partial_{x}u(t,x)-z]}+{\color[rgb]{0,0,0}x^{\prime}(t)\cdot z}$
$=-H(t,x,u(t,x),z)+{\color[rgb]{0,0,0}x^{\prime}(t)\cdot z}$
$=y^{\prime}(t),$
@@
By compactness101010 Note that Assumption 4.2 (vi) prevents a possible blow up of $y^{m}$ as $m\to\infty$. of $\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$ (Remark 4.3), we can, without loss of generality, assume that $(x^{m},y^{m})_{m}$ converges uniformly to a pair $(x^{0},y^{0})\in\mathcal{Y}^{L}(s_{0},x_{0},y_{0},z)$. Let $t\in(s_{0},T]$. Then there is a sequence $(s^{m})_{m}$ in $(s_{0},T]$ with $s^{m}\in\pi^{m}$, $m\in{\mathbb{N}}$, that converges to $t$. Thus $y^{0}(t)=\lim_{m}y^{m}(s^{m})\geq\liminf_{m}u(s^{m},x^{m})\geq u(t,x^{0})$ thanks to the lower semi-continuity of $u$. This concludes the proof. ∎